OpenStudy (anonymous):

find the vertex of f(x) = 3x^2-30x+72 ....I got (5,-51) but the computer keeps saying thats wrong. what did I do wrong?

OpenStudy (anonymous):

x coordinate of vertex = -b/2a = 30/6 = 5 y .............................. = 3*25-150 + 72 = -3 (5,-3)

OpenStudy (anonymous):

your x cood was right - error was in calculating y cood

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

darn! you're right thanks! i couldn't figure it out!

OpenStudy (anonymous):

yw